OpenStudy (atomicreaper):

The difference of a positive integer and its inverse is 15/4. Find the integer.

OpenStudy (atomicreaper):

@welshfella

OpenStudy (welshfella):

the inverse of a positive integer I guess if its ' a' the inverse is '1/a' is that right?

OpenStudy (atomicreaper):

Im not sure, this is the problem.

OpenStudy (welshfella):

its benn a long time..

OpenStudy (welshfella):

yea i think thats right so we have the equation x - 1/x = 15/4 multiply through by 4x 4x^2 - 4 = 15x solve this for x.

OpenStudy (atomicreaper):

4?

OpenStudy (welshfella):

yes - one of the solutions is an integer

OpenStudy (welshfella):

gtg

OpenStudy (atomicreaper):

Ok thanks so much for all the help!!